Solve the initial value problem \(y y^{\prime}+x=\sqrt{x^{2}+y^{2}}\) with \(y(3)=\sqrt{40}\)

a. To solve this, we should use the substitution

\(\boldsymbol{u}=\)


\(u^{\prime}=\)


Enter derivatives using prime notation (e.g., you would enter \(y^{\prime}\) for \(\frac{d y}{d x}\) ).

b. After the substitution from the previous part, we obtain the following linear differential equation in \(\boldsymbol{x}, \boldsymbol{u}, \boldsymbol{u}^{\prime}\)


c. The solution to the original initial value problem is described by the following equation in \(\boldsymbol{x}, \boldsymbol{y}\)



Previous Problem List Next (1 point) Solve the initial value problem yy' + -y2 with y(3) = 40. a. To solve this, we should use the substitution u= help (formulas) help (formulas) u' Enter derivatives using prime notation (e.g., you would enter y' b. After the substitution from the previous part, we obtain the following linear differential equation in x, u, u'. help (equations) c. The solution to the original initial value problem is described by the following equation in c, y. help (equations) Note: You can earn partial credit on this problem.
Answer #1

yy + x = x² + y² Let ut= x² + y² 2u du = 2x +2y du u du at - du =dx (du=fdx to u=x+c Sx²ty2 = x+c Applying the condition y (
Answer #2

IMG_3898.jpg

source: myself
answered by: Gevorg